OpenStudy (mtalhahassan2):

Use the product rule and the power of a function rule to differenite the following functions. Do not simplify. b) (3x^2+4)(3+x^3)^5

OpenStudy (mtalhahassan2):

@Shalante

OpenStudy (anonymous):

This can be solved using the product rule Its just 2 function multiplying together: f(x) and g(x) Lets call \[f(x)=3x^2+4\] \[g(x)=(3+x^3)^5\] Try to find the derivative of both by yourself f'(x)=? g'(x)=? then use this product rule formula f'(x)g(x)+f(x)g'(x)

OpenStudy (mtalhahassan2):

y=(3x^2+4)(3+x^3)^5 =2(3) (3+x^3)^5 + (3x^2+4) 15(3+x^2)

OpenStudy (mtalhahassan2):

is it right???

OpenStudy (jhannybean):

\[f(x) = (3x^2+4)(3+x^3)^5\]\[f(x) = 3x^2+4 ~,~ g(x) = (3+x^2)^5\]

OpenStudy (anonymous):

Derivative of 3x^2 is not 6 Derivative (3+x^3)^5 is not 15(3+x^2)

OpenStudy (jhannybean):

\[\frac{d}{dx} (f(x)) = \frac{d}{dx}(3x^2+4) \cdot (3+x^3)^5 +\frac{d}{dx}((3+x^3)^5) \cdot (3x^2+4)\]

OpenStudy (anonymous):

You forgot the x in the first one You forgot the x^2 in the second one.

OpenStudy (jhannybean):

Remember to apply the chainrule.

OpenStudy (jhannybean):

\[\frac{d}{dx}(f(g(x)) = f'(g(x)) \cdot g'(x)\]

OpenStudy (anonymous):

He did apply the chain rule. He forgot the x on both cases but not the numbers.

OpenStudy (jhannybean):

Ahh ok

OpenStudy (mtalhahassan2):

=2(3x)(3+x^3)^5+(3x^2+4)5(3x^2)(3+x^3)^4 =6x(3+x^3)^5+15x^2(3x^2+4)(3+x^3)^4

OpenStudy (anonymous):

Yes.